# A Typology of Decision-Making Tasks for Visualization

### v1

@@ 7.2.1 Analysis and Reflection of Design
Participants reported that structuring decisions in a graph or hierarchical format using our typology helped them break down problems and decision goals, identify potential decisions that were not initially considered, and compare the current designs with improvements. In particular, the graph served as an externalization of mental models for the participants to better reflect on the overall decision-making process, allowing them to step back from the development of tools and consider the bigger picture. As explained by one of the participants, “I think this [exercise using the typology] was actually very valuable from my standpoint in reflecting on what we are building. Take a step back and reflecting on how these decisions, what we understand, and from an abstraction and typology standpoint, what they mean. So it’s been actually helpful in that regard.”
@@ 7.2.3 Categorization and Organization of Publications
Some participants highlighted the potential of our typology in aiding the research community in categorizing and organizing publications relating to decision-making effectively. In our interviews, three participants utilized our typology to sketch diagrams representing different versions of their tools or extensions published in subsequent papers, demonstrating its capacity to compare decision processes across various tools. Through these diagrams, they were able to discuss the nuances and distinctions between these tools, highlighting areas with varying levels of decision support. This process enabled them to evaluate and compare their own tools’ effectiveness.
The participants expressed that our typology could serve as a valuable tool for evaluating decision support tools due to its ability to organize the decision process into distinct parts and delineate where decisions begin and end. One participant noted, “In my experience as a reviewer, I don’t explicitly ask authors to follow specific patterns or taxonomies to organize their work. However, if they use a taxonomy, it will be easier to follow their paper. Basically, they can organize decision-making papers into these few stages, making it easier to evaluate this type of work.”
@@ 7.3.2 Flexibility to Data and Context
8 Discussion & Future Work
The evaluation of our decision-making task typology received predominantly positive feedback from the participants in our interview study. In this section, we analyze additional feedback from participants to identify potential areas for future research.
8.1 Annotations in the Typology
During the interview study, several participants incorporated annotations into their diagrams or requested such annotations while illustrating their decision-support tools using our typology. These annotations did not expand the set of tasks in our typology or modify their definitions, and they provided additional context for the decisions.
The annotations our participants added included the frequency of decision-making instances, the roles of individuals in a collaborative or organizational setting, and decision-making styles. For instance, one participant highlighted the potential use of our typology to depict decision-making problems based on various “personas,” resulting in distinct diagrams reflecting the roles of decision-makers within an organization. Another participant proposed annotating the diagram to indicate which decision tasks humans perform and which involve computer assistance or collaboration.
Ultimately, the incorporation of annotations highlights an opportunity to further examine decision-making in visualization with our typology. For instance, “personas" could be described through different diagrams for solving the same decision problem. Alternatively, denoting decision tasks performed by humans or computers allows us to concretely express and investigate the role of artificial intelligence in decision-making. Our typology can facilitate deeper exploration and reflection on the many facets that intersect with decision-support visualization systems.
8.2 Measuring the Effectiveness of the Typology
In addition to conducting an interview study with visualization researchers to evaluate our typology, we also assess its effectiveness according to the criteria established by Brehemer et al.[13] and Yi et al.[74]. Specifically, we examine our typology through the lens of three dimensions proposed by Beaudouin-Lafon[10]: (1) descriptive power: “the ability to describe a significant range of existing [designs]”, (2) evaluative power: “the ability to help assess multiple design alternatives”, and (3) generative power: “the ability to help designers create new designs.”
We have showcased the descriptive capabilities of our typology through four case studies and an interview study. In each instance, designers successfully utilized our typology to construct diagrams that effectively described their decision-making problems. Furthermore, we have begun to demonstrate how the typology could facilitate the evaluation and the identification of similar decision problems. Our case studies illustrated the potential of using these diagrams to compare the similarities and differences between domain specific decision problems in a domain-agnostic way.
Finally, we assess the generative aspect of our typology using the definition provided by Brehemer et al.[13]. Specifically, a typology is generative if has a systematic approach for using it to create workflows or diagrams. Our proposed hierarchical decomposition process of decision tasks corresponds with this definition. Starting from the root level, participants demonstrated the ability to recursively expand each node into a composed set of decision tasks, culminating in the creation of a complete diagram. This approach allows visualization designers to methodically externalize decision problems in a structured manner, facilitating a gradual but precise representation of the decision-making process.
8.3 From Typology to Visualization Design
Some participants in our study highlighted the opportunity to integrate the diagrams created using our typology with visualization designs. By pairing these diagrams with their associated visualization designs, it becomes feasible to explore these relationships to develop recommendation systems of decision-support tools, following a similar approach as the Draco system[43]. For instance, the Homefinder Revisited case study, as described in Section 6.1, illustrates a simple diagram representing a multi-criteria decision-making (MCDM) problem. Due to the widespread use of MCDM, this motif occurs in many other, more intricate decision-making scenarios. In the Umbra and Centaurus case studies illustrated in Figures 3 and 4, the same MCDM motif is evident (observed in the top three nodes), indicating that MCDM is a component of both the Umbra and Centaurus decision problems as well. When paired with their associated visualization designs, we can begin to identify and extract visualization design patterns that represent MCDM or other similar sub-decision tasks. In our future work, we aim to investigate how we can integrate our typology diagrams with visualization design processes to assist designers of decision-support tools.
@@ 9 Conclusion
This paper presents a typology for decision-making tasks in visualization, addressing the limitations of existing taxonomies. Built upon prior research and informed by design goals derived from a thorough literature review, the typology comprises three decision tasks: CHOOSE, ACTIVATE, and CREATE. These tasks allow for the representation of complex decision-making structures, as they can be composed or decomposed into other tasks. The typology demonstrates completeness, expressiveness, and utility across various decision-making contexts and is validated through semi-structured interviews with visualization experts.
In summary, this paper contributes potential insights into decision-making in visualization, offering a typology that can be used for analyzing design, facilitating communication, and categorizing visualization publications on decision-making systems. By providing a structured approach to understanding decision-making tasks, the typology offers a tool for researchers and practitioners alike, with implications for improving designs of decision-support tools in the future.

### v3

@@ 7.2.1 Analysis and Reflection of Design
Participants reported that structuring decisions using our typology diagram helped them break down decision problems, identify potential decisions that were not initially considered, and recognize what decisions lacked support in their system. In particular, the typology diagram served as an externalization of mental models for the participants to better reflect on the overall decision-making process, allowing them to step back from the development of tools and consider the bigger picture. As explained by one of the participants, “I think this [exercise using the typology] was actually very valuable from my standpoint in reflecting on what we are building. Take a step back and reflecting on how these decisions, what we understand, and from an abstraction and typology standpoint, what they mean. So it’s been actually helpful in that regard.”
@@ 7.2.3 Categorization and Organization of Publications
Some participants highlighted the potential of our typology in aiding the research community in categorizing and organizing publications relating to decision-making. In our interviews, three participants utilized our typology to sketch diagrams representing different versions of their tools or extensions published in subsequent papers, demonstrating its capacity to compare decision processes across various tools. Through these diagrams, they were able to discuss the nuances and distinctions between these tools, highlighting areas with varying levels of decision support. This process enabled them to evaluate and compare their own tools’ effectiveness.
Some participants expressed that our typology could serve as a valuable tool for evaluating decision support tools due to its ability to organize the decision process into distinct parts and delineate where decisions begin and end. One participant noted, “In my experience as a reviewer, I don’t explicitly ask authors to follow specific patterns or taxonomies to organize their work. However, if they use a taxonomy, it will be easier to follow their paper. Basically, they can organize decision-making papers into these few stages, making it easier to evaluate this type of work.”
@@ 7.3.2 Flexibility to Data and Context
8 Design Implications
Figure 7: Typology diagrams elicited in the first interview. The drawings have been recreated to match the paper’s theme, but the content remains true to what the participant sketched during the interview.
Figure 8: Design improvements and redesigns sketched by participants during the preliminary study. Icons representing visualizations have been added to indicate coordinated views, but these were not part of the participants’ original sketches.
The evaluation presented in Section 7 examines the utility of our typology to describe decision-making problems supported by visualization systems. Here we present initial findings on design implications of our typology gathered from follow-up discussions with available interview participants.
Two of our nine interview participants, a PhD student specializing in design studies (P1) and a visualization developer from industry (P2), were available for a follow-up discussion about their visualization tools. We presented them with the decision-making diagrams they had created during the initial evaluation of our typology. During this session, we asked them to reflect on the diagrams and sketch potential design changes informed by the insights gained. Specifically, they were asked to: 1) identify which decision tasks were well-supported and which lacked adequate support, 2) sketch design improvements for areas where the system did not provide sufficient decision support, and 3) conceptualize a new design from scratch based on the problem outlined in the diagram. Both participants utilized our typology to enhance their designs, incorporating new views, layouts, and interactive features to better support and expand the decisions and processes within their tools.
Using the diagram created in our previous interview, P1 redesigned their tool to better accommodate both novice and expert users. The initial tool design had a workflow where numerous eligibility criteria are evaluated, filtered, and ultimately selected, leading to the creation of more restrictive criteria that feed back into the decision-making process (Figure 7 P1 - W1). This approach catered to novice users lacking the experience necessary to create specific criteria without overlooking crucial variables. However, after creating the decision-making diagram in our first interview, P1 realized that expert users required a different workflow that allows them to create criteria from scratch, exhibited by the second diagram (Figure 7 P1 - W2) drawn in the follow-up discussion. In the initial tool design (Figure 8 P1 - D1), only one of these workflows is supported at a time. In redesigning the tool, the participant integrated both workflows by allowing two separate entry points into the tool (Figure 8 P1 - D2). This redesign supports a broader range of users and allows flexibility in utilizing both workflows when needed. Furthermore, P1 used the decision-making diagrams to reflect further on the design and interaction across the views associated with each decision. This resulted in P1 adding coordinated views to facilitate the ACTIVATE and CREATE decisions, and replacing an overview visualization with more interactive capabilities to support experts in the CREATE decision.
P2 reflected on the lack of support in their visualization tool for some of the decisions contained in their decision diagram. The tool primarily supports the exploratory data analysis stage (EDA WORKFLOW (*) in Figure 7 P2 - W) where users iteratively explore data and make intertwined decisions, followed by a sequential stage that culminates in a “CREATE report” decision (the CHOOSE $\rightarrow$ CREATE motif in Figure 7 P2 - W). P2 noted that the final CREATE decision is not well-supported by the tool, primarily because the initial requirements gathering focused more on the iterative, exploratory, and data-driven aspects of the process, rather than the sequential stages of decision-making. As a result, tasks like creating presentation slides are left to existing tools rather than being integrated into the system. However, P2 noted that a possible improvement in their tool that allows users to save intermediate results during exploration could better support the final CREATE decision. To address this, P2 sketched and considered two possible changes in the tool design: a “shoebox” view where users drag and drop elements and insights gathered during the EDA process (Figure 8 P2 - D1), and a split view with the ability to drag and drop results directly into presentation slides (Figure 8 P2 - D2)). P2 ultimately preferred the split view design as it provides continuous visibility and directly supports the final CREATE decision.
While these are only preliminary findings, there is evidence that our typology aids in the design of visualization systems. In future work, we intend to investigate further how our typology assists in the design of decision-support visualization systems.
9 Limitations
A few limitations of our typology emerged during our interview study: (1) the imprecise description of decision-making context, (2) the flexibility in diagram decomposition, and (3) the lack of guidance for final design decisions.
First, the typology describes decision-making tasks, but it does not account for broader contexts in which these tasks can occur. During the interview study, several participants supplemented their diagrams with annotations to provide context, without altering the core tasks or their definitions. These annotations included details such as the frequency of decision-making instances, roles within collaborative or organizational settings, and decision-making styles. While these annotations do not change the core definitions of the tasks, they often influence the design of decision-support systems. For example, one participant suggested using the typology to depict decision-making based on different "personas," resulting in diagrams that reflect distinct roles within an organization, while another participant proposed annotating diagrams to differentiate between tasks performed by humans and those involving computer assistance or collaboration.
Second, the typology’s flexibility in decomposing decision-making tasks also presents a challenge. There are no rules or exact guidance on the appropriate level of granularity to represent decision-making problems. Our participants annotated the inputs and outputs of each decision task on the arrows that connect them. However, across participants, we have seen that the level of detail varied significantly–some provided high-level information, while others included very specific dataset details. This lack of a stopping point in the recursive decomposability of the tasks mirrors limitations seen in other typologies, such as the under specified abstraction levels found in Brehmer et al.’s typology [13]. While this flexibility can ease diagram creation, it presents challenges in maintaining consistency.
Lastly, we primarily focused on the descriptiveness of decision-making problems, and how to use these descriptions as tools in visualization design. However, our typology offers no direct mechanism to compare designs that satisfy the same decision-making diagram. We acknowledge that selecting between design alternatives is influenced by various factors beyond the typology itself, such as the designer’s experience, time constraints, and the decision-making style of the domain expert, which are not captured by our typology.
While our typology offers valuable insights and flexibility in describing decision-making processes, these limitations indicate areas for refinement and further exploration. In the following section, we outline our future work, which will address some of these limitations and expand upon our typology’s potential applications.
10 Future Work
While the limitations section outlined some constraints of our typology, future work will focus on leveraging the typology as a foundation for building new visualization knowledge, applications, and tools. The following points summarize our intended directions:
Full-Scale Design Evaluation. A comprehensive design evaluation with designers and their domain experts will be conducted to explore the typology’s utility in end-to-end visualization design settings. This will involve longitudinal studies rather than brief interviews, allowing for deeper insights into how the typology can be applied in real-world design processes.
Expanding Design Integration. Informed by our user studies, future work will involve integrating the diagrams generated by our typology with corresponding visualization designs, potentially leading to the development of recommendation systems for decision-support tools, such as the Draco system [46]. We plan to conduct an in-depth analysis of recurring decision-making patterns, such as multi-criteria decision-making, which could evolve into an open-source project where contributors share typology diagrams. This would enable the study of decision-making patterns and the recommendation of similar problems and design solutions.
Connecting Decision-Making and Analytical Tasks. Another avenue for future work is exploring connections between decision-making tasks in our typology and established analytical tasks, such as those in Amar and Eagan’s taxonomy [3] or Brehmer et al.’s typology [13]. Identifying patterns in these connections could help predict decision-making tasks based on user interaction data or suggest relevant analytical tasks based on decision goals. This would have a direct impact on the reasoning about decision-making tasks in the context of visual analytics, as it can enhance the alignment between decision support systems and the analytical workflows that users engage with, ultimately improving the efficiency of visual analytics tools.
Human-Machine Decision-Making. We plan to use our typology to explore and reason about the division between human and machine decision-making in systems, and how designs can reflect or leverage this distinction. We believe our typology can provide insights into human-machine collaboration, the effects of automating parts of the decision-making process, and the role of human intervention in AI-driven systems. This is especially crucial in visual analytics, where effective collaboration between humans and AI ensures interpretability, trust, and control over complex decision-support tools.
Overall, the next steps will involve applying our typology to the development and improvement of tools and decision-making processes, seeking to validate its relevance and effectiveness in real-world design and decision-making scenarios.
@@ 11 Conclusion
This paper presents a typology for decision-making tasks in visualization, addressing the limitations of existing task taxonomies. Built upon prior research and informed by design goals derived from a thorough literature review, the typology comprises three decision tasks: CHOOSE, ACTIVATE, and CREATE. These tasks allow for the representation of complex decision-making structures, as they can be composed or decomposed into other tasks. The typology demonstrates completeness, expressiveness, and utility across various decision-making contexts and is validated through semi-structured interviews with visualization experts. During the interviews, the visualization experts expressed that the typology has the potential to be used in analyzing designs of decision-support visualization systems, facilitating communication between designers and domain users, and categorizing visualization publications on decision-making. By providing a structured approach to understanding decision-making tasks, the typology offers a tool for researchers and practitioners alike, with implications for improving designs of decision-support tools in the future.
Appendix A Supplemental Materials
A.1 Quotes Supporting the Design Goals for a Decision Task Typology
  1. G1
    • By working with the domain experts, we constructed a three-stage iterative workflow for analyzing and improving bus networks[75]
  2. G2
  3. G3
  4. G4
    Information Processing-Focused: The decision tasks within our typology should demonstrate flexibility and impartiality towards input and output types, instead emphasizing how information is processed and evaluated.
A.2 Diagrams Created by Our Participants
Figure 10: Diagrams created by our participants during the interview study. Texts, labels, and annotations that contain identifiable information have been removed to preserve anonymity. We labeled the diagrams with the domain of application.